OpenStudy (anonymous):

expand each logarithm: 1. log x^3 y^5 2.log[base 7] 22xyz 3. log[base 4] 5 [square root x] 4.log 3m^4 n^-2 5. log [base 5] r/s 6. log [ base 3] (2x)^2

OpenStudy (unklerhaukus):

\[\log(ab)=\log(a)+\log(b)\] \[\log(c^n)=n\log c\]

OpenStudy (anonymous):

so wod the first answer be x log 3 + y log 5? or log x ^3 + log x^5

OpenStudy (unklerhaukus):

\[\log x^3 y^5=\log x^3+\log y^5=\]

OpenStudy (anonymous):

3 log x + 5 log y ?

OpenStudy (unklerhaukus):

thats right!

OpenStudy (anonymous):

=D so how about the second one ?

OpenStudy (anonymous):

wod it be 22 log [base 7] x + 22 log [base 7} y + 22 log [base 7} z ??

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[\log_7 22xyz=\log_722+\log_7x+\log_7y+\log_7z\]

OpenStudy (anonymous):

okay how about number 3.. i dont know how to solve it ?

OpenStudy (unklerhaukus):

\[\log_4 5 \sqrt x=\]

OpenStudy (unklerhaukus):

break the log into a sum of logs, then use \(\sqrt x=x^{1/2}\)

OpenStudy (anonymous):

okay so would it be : 1/2 log [base of 4 ] 5

OpenStudy (unklerhaukus):

not quite, your answer should have a sum of logs in it

OpenStudy (anonymous):

like wat do u mean ? like 5 log [base of 4 ] + log 1/2

OpenStudy (unklerhaukus):

thats closer

OpenStudy (anonymous):

log [base of 4] 5 + log 1/2 ?

OpenStudy (unklerhaukus):

thats is eve closer , but where did x go?

OpenStudy (anonymous):

oh log [base of 4 ] 5 + log 1/2x

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

\[\log_4 5 \sqrt x=\log_4 5 +\log_4\sqrt x\]\[\qquad\qquad=\log_45+\log_4x^{1/2}\]\[\qquad\qquad=\log_45+\frac12\log_4x\]

OpenStudy (anonymous):

oh okay i get it now , thank you :) so th

OpenStudy (anonymous):

so then number 4 wod be: log 3 + 4 log m + -2 log n ? is that right??

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

you've got it!

OpenStudy (unklerhaukus):

for 5. you have to remember that \[\frac 1s=s^{-1}\]

OpenStudy (anonymous):

so it would be : log [base of 5] 1 + -1 log s ???

OpenStudy (unklerhaukus):

where did r go?

OpenStudy (anonymous):

1 log [base of 5] r + -1 log [base of 5 ] s ????

OpenStudy (unklerhaukus):

thats good , but you dont need the ones (they are implied) and +- can be written as just -

OpenStudy (anonymous):

oh okay so it is : log [base of 5] r - log [base of 5 ] s ??

OpenStudy (unklerhaukus):

that is right .

OpenStudy (anonymous):

okay thank you and then number 6 wod be : 2 log [base of 3] 2x + 2 log [base of 3] 2x ??? is that right ?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

check that one again

OpenStudy (anonymous):

would it be : 2 log [base of 3] 2 + 2 log [base of 3] x

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

\[\LARGE\color{red}\checkmark\]

OpenStudy (anonymous):

thank you so much for ur help :)

OpenStudy (anonymous):

@UnkleRhaukus okay i have one more question : wat wod be the answer for log[base 3] 7 ( 2x - 3) ^2

OpenStudy (unklerhaukus):

you tell me

OpenStudy (anonymous):

wod the answer be : 2 log[base 3 ] 7 + 2 log[base 3] 2 + 2 log [base 3 ] x + 2 log [base 3] -3

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

not quite

OpenStudy (unklerhaukus):

all you can do is this\[\log(a(b+c)^n)=\log(a)+\log(b+c)^n\]\[\qquad\qquad\qquad=\log(a)+n\log(b+c)\]

OpenStudy (unklerhaukus):

the log a sum (or difference) cannot be simplified

OpenStudy (anonymous):

so it would be log [base 3] 7 + 2 log [base 3] ( 2x -3 ) ?

OpenStudy (unklerhaukus):

yeah that is all that can be done , great work

OpenStudy (anonymous):

=D